A tennis ball is dropped from a certain height it falls on the floor and bounces back. Assuming that the upward motion of velocity is positive what will be the graph of the situation? 

Grade:11

1 Answers

Vikas TU
7 years ago
when it id dropped,
-h = 0 – 0.5gt^2
t = root(2h/g) ..............(1)
final velocity becomes,
v = root(2gh)..........(2)
divide the eqns. (!) and  (2)
v/t = g =>  v = gt which is linear. for time t.
After bouncing,
v’ = (2gh) – 2gh’
and
h’ = 2ght’ – 0.5gt’^2
or
v’ = 2gh – 2g( 2ght’ – 0.5gt’^2) which is parabolic for time t’.
Hence the graph would be linear for interval time t aand then after it would become parabolic for time interval t’.